OpenStudy (anonymous):

Suppose you have a function f(x) = e^(-Ax), where A is a positive real number. Show that the integral from 1 to 2 of x*f(x) dx approaches 0 as A approaches infinity

OpenStudy (anonymous):

Otherwise: \[f(x) = e ^{-Ax}\], A is a positive real number Show: \[\int\limits_{1}^{2}xf(x)dx\] approaches 0 as A approaches infinity

OpenStudy (anonymous):

Try to consider that \[\int\limits_{}^{}e^{nx}dx=\frac{ e^{nx} }{n}\] As n approaches negative infinity, you get \[\frac{ e^{-\infty x} }{-\infty}=-\frac{1}{e^{\infty x} \infty}\] Which for x = 1 and x = 2 approaches 0

OpenStudy (anonymous):

That was actually the first part of the problem. Wouldn't integrating by x times the function be more different that just that answer, even though in the end it will come down to a bunch of fractions that will ultimately reach zero as a goes to infinity?

OpenStudy (anonymous):

Oh, didn't read that x, sorry. I'll just post the WolframAlpha approach, since I have no idea on how to integrate by parts. http://www3.wolframalpha.com/Calculate/MSP/MSP76951a39f64hh908gi9d000029a17226icg50i1h?MSPStoreType=image/png&s=1&w= [[1]]&h=[[2]] From that it should be easy.

OpenStudy (anonymous):

(You have to copy-paste the full URL into your browser, not just the part in blue)

OpenStudy (anonymous):

Ah. Thank you for the that. I actually integrated that using integration by parts and that just confirmed my answer. Thanks!

OpenStudy (anonymous):

If that's all you needed you can try to go on WolframAlpha.com next time.